OpenStudy (anonymous):

Solve 25x= 125x - 7 A) -7 B) 21 C) 21/7

OpenStudy (anonymous):

You already have the answers, have you tried plugging them in? Or better yet, trying at all?

OpenStudy (zepp):

25x = 125x - 7 Put the x terms together 125x -25x = 7 100x = 7

OpenStudy (anonymous):

25(-7)=125(-7)-7 -175 doesnt equal -882 25(21)=125(21)-7 525 doesnt equal 2618 25(21/7)=125(21/7)-7 75 doesnt equal 368 um. either my calculator is broken or you need another answer. i got x= 7/100 :|

OpenStudy (zepp):

Pretty sure that 7/100 is D) :D

OpenStudy (anonymous):

yeah..

OpenStudy (anonymous):

nope d was wrong

OpenStudy (anonymous):

0.07? was that a choice?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the only right choices are the ones left. so i just guesesed. thanksk though